# Swarm-based gradient descent meets simulated annealing

### v1

@@ 5.2. Large time behavior
Putting these two lemmas together, we can prove the theorem.
Proof for (2.7) in Theorem 2.2.
We deploy the proof by contradiction. This is to assume (2.7) is not true, meaning $\liminf_{t\rightarrow\infty}\overline{\mathscr{F}}^{\mu}_{t}>F_{*}$. Then according to Lemma 5.3, there exists $\epsilon_{0}>0$ such that ¯$\overline{\mathscr{F}}^{\mu}_{t}>F_{*}+\epsilon_{0}$ for all $t\geqslant 0$. Setting $\epsilon=\epsilon_{0}$ in Lemma 5.4, mass is not conserved, contradicting the property of the equation. Hence (2.7) has to be true, namely:
@@
We now spell out details for the two lemmas.
Proof for Lemma 5.3.
This is also completed by contradiction. Assuming the statement is not true, meaning there is no such $\epsilon_{0}$ to make ¯$\overline{\mathscr{F}}^{\mu}_{t}>F_{*}+\epsilon_{0}$, then $\inf_{t\geqslant 0}\overline{\mathscr{F}}^{\mu}_{t}=F_{*}$. Considering $\liminf_{t\rightarrow\infty}\overline{\mathscr{F}}^{\mu}_{t}>F_{*}$, there must exist a finite time, denoted as $t_{*}\geqslant 0$ such that ¯$\overline{\mathscr{F}}^{\mu}_{t_{*}}=F_{*}$. According to Assumption 2.2, $F$ has a unique global minimum at $\mathbf{x}_{*}$, so at $t_{*}$, $\mu_{t_{*}}(\mathbf{x},m)$ must have the form of
(5.12) $\mu_{t_{*}}(\mathbf{x},m)=\delta_{\mathbf{x}_{*}}(\mathbf{x})\otimes\eta(m)\,.$
We will show below that $\mu_{t_{*}}(\mathbf{x},m)$ in the form of (5.12) is an equilibrium. If so, $\mu_{t}(\mathbf{x},m)=\mu_{t_{*}}(\mathbf{x},m)$ for all $t>t_{*}$, and thus ¯$\overline{\mathscr{F}}^{\mu}_{t}=F_{*}$ for all time after $t_{*}$. This contradicts the assumption that $\liminf_{t\rightarrow\infty}\overline{\mathscr{F}}^{\mu}_{t}>F_{*}$, completing the proof.
To show $\mu_{t_{*}}(\mathbf{x},m)$ in the form of (5.12) is an equilibrium is amount to plugging the form into the equation (2.4a) and show the right hand side vanishes. Since this is a probability-measured solution with a Dirac delta on $\mathbf{x}$-domain, the proof has to be conducted in the weak sense. Specifically, given any smooth test function $\phi(\mathbf{x},m)$, we show the right hand side of (2.4a), when tested by $\phi(\mathbf{x},m)$, vanishes:
[…]
We will show all three terms are zero. Indeed,
  • Since $\mu_{t}$ takes on the form of (5.12),
    (5.14) $\mathrm{Term\ I}=$ $\iint(F(\mathbf{x})-\overline{\mathscr{F}}^{\mu}_{t}m\partial_{m}\phi(\mathbf{x},m)\hskip 0.56917pt\textnormal{d}\mu_{t}(\mathbf{x},m)$
  • Similarly,
    (5.15) $\mathrm{Term\ II}=$ $\iint\nabla_{\mathbf{x}}\phi(\mathbf{x},m)\cdot\nabla F(\mathbf{x})\delta_{x_{*}}(\mathbf{x})\textnormal{d}{\mathbf{x}}\textnormal{d}\eta(m)$
  • To show $\mathrm{Term\ III}=0$, we note that
    (5.16) $\mathrm{Term\ III}=$ $\iint\sigma(m)\Delta_{\mathbf{x}}\phi(\mathbf{x},m)\delta_{x_{*}}(\mathbf{x})\textnormal{d}{\mathbf{x}}\textnormal{d}\eta(m)$
Plugging (5.14)-(5.16) into (5.13), it is clear that $\mu_{t_{*}}(\mathbf{x},m)=\delta_{\mathbf{x}_{*}}(\mathbf{x})\otimes\eta(m)$ is an equilibrium, and thus $\mu_{t}(\mathbf{x},m)=\mu_{t_{*}}(\mathbf{x},m)$ for all $t>t_{*}$, making ¯$\overline{\mathscr{F}}^{\mu}_{t}=F_{*}$ and contradicting the assumption that $\liminf_{t\rightarrow\infty}\overline{\mathscr{F}}^{\mu}_{t}>F_{*}$. ∎
Proof for Lemma 5.4.
  1. (a)
  2. (b)
Indeed, we recall its trajectory governed by the SDE:
$m_{s}(t)>\exp\left(\frac{\epsilon}{2}(t-T_{c})\right)m_{s}(T_{c})>m_{c}\,,$
confirming (a)-(b). Accordingly,
(5.17) $\int\limits_{{\mathcal{N}}_{\epsilon/2}}$ $\int\limits^{\infty}_{m_{c}}m\hskip 0.56917pt\textnormal{d}\mu_{t}(\mathbf{x},m)=\mathbb{E}(m(t)|\mathbf{x}(t)\in{\mathcal{N}}_{\epsilon/2},m(t)>m_{c})\mathbb{P}\left(\left\{\mathbf{x}(t)\in{\mathcal{N}}_{\epsilon/2},m(t)>m_{c}\right\}\right)$
[…]
completing the proof. ∎
Proof for (2.10).
We again deploy a proof by contradiction and the argument is rather similar to that of (2.7). Assuming that it is not true, then for all $t\leqslant t_{\epsilon}(m_{c})$,
@@
Since the probability is non-trivial, one can sample a point in the set of ${\mathcal{N}}_{\epsilon}\times(m_{c},\infty)$. Starting with $\mathbf{x}(\eta_{\epsilon})\in{\mathcal{N}}_{\epsilon}$ and $m(\eta_{\epsilon})>m_{c}$, with the same calculation as in (5.17), we get:
@@ 6. Numerical experiments
For the numerical evidence we will use the following $\sigma$ function in the application of Algorithm 1:
@@
The smooth decay of the function (see Figure 2(a)) gives different magnitudes of noise to each particle depending on their mass, but eventually for sufficiently high mass the noise will disappear.
@@ 6.1.1. Ackley function in lower dimensions
Figure 6.5. Swarm movement (in black) for Ackley function in $d=2$.
@@ 6.1.3. Ackley function in higher dimensions
The fourth experiment is to evaluate the performance of the algorithm on a high dimensional function (Ackley function for $d=10$). The hope is the Brownian motion in the dynamics can help overcoming the curse of dimensionality. We set the parameters to be $\lambda=1$, $\beta=1/N$. In Figure 6.9, we show the convergence in time of $20$ runs of experiments using Algorithm 1, and compare it to the solution to the deterministic system shown in (4.1). The observation is that the deterministic system converges very quickly at the initial steps, but it saturates at local minima, and the global optimization is never achieved. On the contrary, Algorithm 1, using either $2000$ or $200$ samples both achieve the global minimum, albeit with slower convergence rate. We attribute this to the randomness in the method and this observation confirms Theorem 4.1. We also emphasize that the performance of the algorithm using smaller number of samples ($200$) is compatible to that using $2000$, indicating the overcome of dimensionality.
@@ 6.2. Numerical study
with parameters chosen to be $d=1$ and $D=0$, and the same choices of $\{x_{i}\}$. In this case, we run the simulation for a very long time till $t\sim 700$. As seen in Figure 6.11, the error seems to saturate at around $O(1)$ after $t=10$, and the second dip occurs at $t=690$, dragging the error down to 00. The convergence takes place in a staggered pattern. The conjecture is that the value of the local minimum of the Rastrigin function is very close to that of the global one, and the samples stuck at the local minimum for a very long stretch of time, accumulating all the mass. The lighter “explorer" uses this time to explore and find the global minimum. Before the very small basin gets found, local minimum takes the lead and the error stays at $O(1)$, and the second dip occurs when the global basin finally was found by one explorer, and the mass gets transferred to it. We find this staggered pattern enlightening, and future research is needed for spelling out the explicit dependence of this behavior on different parameters.
@@
(A.6) $\begin{split}\frac{\textnormal{d}}{\textnormal{d}t}\left(\sum_{j=1}^{N}\right.&\left.m^{j}(t)\phi(\mathbf{x}^{j}(t))\right)=\\ &-\sum_{j=1}^{N}m^{j}_{t}\underbrace{\nabla_{\mathbf{x}}\phi(\mathbf{x}^{j}(t))\cdot\nabla F(\mathbf{x}^{j}(t))}_{\dot{\mathbf{x}}=-\nabla F(\mathbf{x}(t))}-\sum_{j=1}^{N}\phi(\mathbf{x}^{j}(t))\underbrace{m^{j}(t)(F(\mathbf{x}^{j}(t))-\overline{F}(t))}_{\dot{m}=-m(t)(F(\mathbf{x}(t)-\overline{\mathscr{F}}^{\nu}(t))},\end{split}$
@@ Appendix B Technical lemmas
The following lemma is deployed in the proof for the long time behavior Lemma 5.4.
@@
Under the condition of Theorem 2.2. Let $\mu_{t}(\mathbf{x},m)$ solve (2.4a). If at a time $t_{*}$ so that $\mu_{t_{*}}(\mathbf{x},m)$ takes on the form of $\delta_{\mathbf{x}_{*}}\otimes\eta(m)$ for some probability measure $\eta$, then $\mathrm{Supp}_{m}(\cdot)\subset[m_{c},\infty)$.
@@
where the $\big{(}F(\mathbf{x})-\overline{\mathscr{F}}^{\mu}_{t}\big{)}\partial_{m}(m\mu)\left.\right|_{\mu_{t_{*}}(\mathbf{x},m)}=0$ and is dropped out upon testing, using the same argument that showed (5.14). We will show that both $\mathrm{Term\ I}$ and $\mathrm{Term\ II}$ are zero.

### v2

@@ 5.2. Large time behavior
The proof of these two lemmas are found in Appendix B [^6] The two lemmas require the assumption of a unique global minimum as stated in Assumption 2.2. However, in the proof in Appendix B, we will discuss that this assumption can be relaxed, but analysis needs to be altered.. Putting these two lemmas together, we can prove the theorem.
Proof of Theorem 2.2.
We prove (2.7) using proof by contradiction. This is to assume (2.7) is not true, meaning $\liminf_{t\rightarrow\infty}\overline{\mathscr{F}}^{\mu}_{t}>F_{*}$. Then according to Lemma 5.3, there exists $\epsilon_{0}>0$ such that ¯$\overline{\mathscr{F}}^{\mu}_{t}>F_{*}+\epsilon_{0}$ for all $t\geqslant 0$. Setting $\epsilon=\epsilon_{0}$ in Lemma 5.4, mass is not conserved, contradicting the property of the equation. Hence (2.7) has to be true, namely:
@@
To show (2.10), we again deploy a proof by contradiction and the argument is rather similar to that of (2.7). Assuming that it is not true, then for all $t\leqslant t_{\epsilon}(m_{c})$,
@@
Since the probability is non-trivial, one can sample a point in the set of ${\mathcal{N}}_{\epsilon}\times(m_{c},\infty)$. Starting with $\mathbf{x}(\eta_{\epsilon})\in{\mathcal{N}}_{\epsilon}$ and $m(\eta_{\epsilon})>m_{c}$, with the same calculation as in (B.6), we get:
@@ 6. Numerical experiments
For the numerical evidence in Section 6.1 we will use the following $\sigma$ function in the application of Algorithm 1:
@@
The smooth decay of the function (see Figure 2(a)) gives different magnitudes of noise to each particle depending on their mass, but eventually for sufficiently high mass the noise will disappear. We only present three examples in this section. Extensive numerical experiments are presented in Appendix C.
@@ 6.1.1. Ackley function in lower dimensions
Figure 6.5. Swarm movement (in red) for Ackley function in $d=2$.
@@ 6.1.3. Ackley function in higher dimensions
The fourth experiment is to evaluate the performance of the algorithm on a high dimensional function (Ackley function for $d=10$). The hope is the Brownian motion in the dynamics can help overcoming the curse of dimensionality. We set the parameters to be $\lambda=1$, $\beta=1/N$. In Figure 6.9, we show the convergence in time of $20$ runs of experiments using Algorithm 1, and compare it to the solution to the deterministic system shown in (4.1). The observation is that the deterministic system converges very quickly at the initial steps, but it saturates at local minima, and the global optimization is never achieved. On the contrary, Algorithm 1, using either $2000$ or $200$ samples both achieve the global minimum, albeit with slower convergence rate. We attribute this to the randomness in the method and this observation confirms Theorem 4.1. We also emphasize that the performance of the algorithm using smaller number of samples ($200$) is compatible to that using $2000$, so a small number of samples already have demonstrated impressive numerical performance in this example.
@@ 6.2. Numerical study
with parameters chosen to be $d=1$ and $D=0$, and the same choices of $\{x_{i}\}$. In this case, we run the simulation for a very long time till $t\sim 700$. As seen in Figure 6.11, the error seems to saturate at around $O(1)$ after $t=10$, and the second dip occurs at $t=690$, dragging the error down to 00. The convergence takes place in a staggered pattern. The conjecture is that the value of the local minimum of the Rastrigin function is very close to that of the global one, and the samples stuck at the local minimum for a very long stretch of time, accumulating all the mass. The lighter “explorer” uses this time to explore and find the global minimum. Before the very small basin gets found, local minimum takes the lead and the error stays at $O(1)$, and the second dip occurs when the global basin finally was found by one explorer, and the mass gets transferred to it. We find this staggered pattern enlightening, and future research is needed for spelling out the explicit dependence of this behavior on different parameters.
@@
Acknowledgments
ET thanks the Fondations Sciences Mathématiques des Paris (FSMP) and LJLL at the Sorbonne University for the support and hospitality. All four authors thank the two anonymous reviewers for constructive and insightful suggestions.
@@ Acknowledgments
(A.6) $\begin{split}\frac{\textnormal{d}}{\textnormal{d}t}\Bigg{(}\sum_{j=1}^{N}&m^{j}(t)\phi(\mathbf{x}^{j}(t))\Bigg{)}=\\ &-\sum_{j=1}^{N}m^{j}_{t}\underbrace{\nabla_{\mathbf{x}}\phi(\mathbf{x}^{j}(t))\cdot\nabla F(\mathbf{x}^{j}(t))}_{\dot{\mathbf{x}}=-\nabla F(\mathbf{x}(t))}-\sum_{j=1}^{N}\phi(\mathbf{x}^{j}(t))\underbrace{m^{j}(t)(F(\mathbf{x}^{j}(t))-\overline{F}(t))}_{\dot{m}=-m(t)(F(\mathbf{x}(t)-\overline{\mathscr{F}}^{\nu}(t))},\end{split}$
@@ Appendix B Technical lemmas
We provide proofs of Lemma 5.3-Lemma 5.4, and Lemma B.1 that was deployed wherein.
Proof of Lemma 5.3.
This is obtained by contradiction. Assuming the statement is not true, meaning there is no such $\epsilon_{0}$ to make ¯$\overline{\mathscr{F}}^{\mu}_{t}>F_{*}+\epsilon_{0}$, then $\inf_{t\geqslant 0}\overline{\mathscr{F}}^{\mu}_{t}=F_{*}$. Considering $\liminf_{t\rightarrow\infty}\overline{\mathscr{F}}^{\mu}_{t}>F_{*}$, there must exist a finite time, denoted as $t_{*}\geqslant 0$ such that ¯$\overline{\mathscr{F}}^{\mu}_{t_{*}}=F_{*}$. According to Assumption 2.2, $F$ has a unique global minimum at $\mathbf{x}_{*}$, so at $t_{*}$, $\mu_{t_{*}}(\mathbf{x},m)$ must have the form of [^8] We note here that if there are multiple global minima $\{\mathbf{x}_{k}\}_{k=1}^{K}$, the form of $\mu_{t_{*}}(\mathbf{x},m)$ becomes $\mu_{t_{*}}(\mathbf{x},m)=\sum_{k=1}^{K}\delta_{\mathbf{x}_{k}}(\mathbf{x})\otimes\eta_{k}(m)$ where $\sum_{k}\eta_{k}$ has measure $1$. The rest of the proof stays the same, with $\mathbf{x}_{*}$ replaced by the collection of $\mathbf{x}_{k}$. The proof of this lemma depends on Lemma B.1, which can also be revised to accommodate the multiple global minima case.
(B.1) $\mu_{t_{*}}(\mathbf{x},m)=\delta_{\mathbf{x}_{*}}(\mathbf{x})\otimes\eta(m)\,.$
We will show below that $\mu_{t_{*}}(\mathbf{x},m)$ in the form of (B.1) is an equilibrium. If so, $\mu_{t}(\mathbf{x},m)=\mu_{t_{*}}(\mathbf{x},m)$ for all $t>t_{*}$, and thus ¯$\overline{\mathscr{F}}^{\mu}_{t}=F_{*}$ for all time after $t_{*}$. This contradicts the assumption that $\liminf_{t\rightarrow\infty}\overline{\mathscr{F}}^{\mu}_{t}>F_{*}$, completing the proof.
To show $\mu_{t_{*}}(\mathbf{x},m)$ in the form of (B.1) is an equilibrium amounts to plugging the form into the equation (2.4a) and show the right hand side vanishes. Since this is a probability-measured solution with a Dirac delta on $\mathbf{x}$-domain, the proof has to be conducted in the weak sense. Specifically, given any smooth test function $\phi(\mathbf{x},m)$, we show the right hand side of (2.4a), when tested by $\phi(\mathbf{x},m)$, vanishes:
[…]
We will show all three terms are zero. Indeed,
  • Since $\mu_{t}$ takes on the form of (B.1),
    (B.3) $\mathrm{Term\ I}=$ $\iint(F(\mathbf{x})-\overline{\mathscr{F}}^{\mu}_{t})m\partial_{m}\phi(\mathbf{x},m)\hskip 0.56917pt\textnormal{d}\mu_{t}(\mathbf{x},m)$
  • Similarly,
    (B.4) $\mathrm{Term\ II}=$ $\iint\nabla_{\mathbf{x}}\phi(\mathbf{x},m)\cdot\nabla F(\mathbf{x})\delta_{x_{*}}(\mathbf{x})\textnormal{d}{\mathbf{x}}\textnormal{d}\eta(m)$
  • To show $\mathrm{Term\ III}=0$, we note that
    (B.5) $\mathrm{Term\ III}=$ $\iint\sigma(m)\Delta_{\mathbf{x}}\phi(\mathbf{x},m)\delta_{x_{*}}(\mathbf{x})\textnormal{d}{\mathbf{x}}\textnormal{d}\eta(m)$
Plugging (B.3)-(B.5) into (B.2), it is clear that $\mu_{t_{*}}(\mathbf{x},m)=\delta_{\mathbf{x}_{*}}(\mathbf{x})\otimes\eta(m)$ is an equilibrium, and thus $\mu_{t}(\mathbf{x},m)=\mu_{t_{*}}(\mathbf{x},m)$ for all $t>t_{*}$, making ¯$\overline{\mathscr{F}}^{\mu}_{t}=F_{*}$ and contradicting the assumption that $\liminf_{t\rightarrow\infty}\overline{\mathscr{F}}^{\mu}_{t}>F_{*}$. ∎
Proof of Lemma 5.4.
  1. (a)
  2. (b)
Indeed, we recall its trajectory governed by the SDE:
$m_{s}(t)>\exp\left(\frac{\epsilon}{2}(t-T_{c})\right)m_{s}(T_{c})>m_{c}\,,$
confirming (a)-(b). Accordingly,
(B.6) $\int\limits_{{\mathcal{N}}_{\epsilon/2}}$ $\int\limits^{\infty}_{m_{c}}m\hskip 0.56917pt\textnormal{d}\mu_{t}(\mathbf{x},m)=\mathbb{E}(m(t)|\mathbf{x}(t)\in{\mathcal{N}}_{\epsilon/2},m(t)>m_{c})\mathbb{P}\left(\left\{\mathbf{x}(t)\in{\mathcal{N}}_{\epsilon/2},m(t)>m_{c}\right\}\right)$
[…]
completing the proof. ∎
@@
Under the condition of Theorem 2.2. Let $\mu_{t}(\mathbf{x},m)$ solve (2.4a). If at a time $t_{*}$ so that $\mu_{t_{*}}(\mathbf{x},m)$ takes on the form of $\delta_{\mathbf{x}_{*}}\otimes\eta(m)$ for some probability measure $\eta$, then $\mathrm{Supp}(\eta(m))\subset[m_{c},\infty)$. [^9] We note that in the statement of the lemma, we assume $\mu_{t_{*}}(\mathbf{x},m)$ has a form of one Dirac delta located on $\mathbf{x}_{*}$. In the situation when there are multiple global minima $\{\mathbf{x}_{k}\}_{k=1}^{K}$, the proof can be adjusted to handle the form of $\mu_{t_{*}}(\mathbf{x},m)=\sum_{k=1}^{K}\delta_{\mathbf{x}_{k}}\otimes\eta_{k}(m)$, with the conclusion being $\mathrm{Supp}(\eta_{k}(m))\subset[m_{c},\infty)$ for all $k$. In the proof, we should employ a list of test functions with each taking the form of $\phi_{k}(\mathbf{x},m)=|\mathbf{x}-\mathbf{x}_{k}|^{2}\xi_{k}(\mathbf{x})$.
@@
where the $\big{(}F(\mathbf{x})-\overline{\mathscr{F}}^{\mu}_{t}\big{)}\partial_{m}(m\mu)\left.\right|_{\mu_{t_{*}}(\mathbf{x},m)}=0$ and is dropped out upon testing, using the same argument that showed (B.3). We will show that both $\mathrm{Term\ I}$ and $\mathrm{Term\ II}$ are zero.
@@
Appendix C Numerical Experiments
In this section we present further numerical evidence of the performance of our Swarm-based Simulated Annealing (SSA) Algorithm 1 for different types of 2D non-convex functions. We used (6.1) as our $\sigma$ function with different parameters and different number of particles in each case. These examples are selected from [27], and in each plot below, we show the landscape of the objective function, the convergence of the provisional minimum in time, and four time snapshots of available samples. At initial time, the samples are purposely removed from the global basin.
Figure C.12. SSA applied to a Levy function: $F(x,y)=\sin^{2}(\pi w(x))+(w(x)-1)^{2}[1+10\sin^{2}(\pi w(x)+1)]+(w(y)-1)^{2}[1+\sin^{2}(2\pi w(y))]$, with $w(x)=1+\frac{x-1}{4}$. The parameters used are $N=16$, $\lambda=2$ and $\beta=1/8$. The global minimum is $F(x_{\ast},y_{\ast})=0$ at $(x_{\ast},y_{\ast})=(1,1)$. Top Left: landscape of he function. Lower Left: long time behavior. Center and Right: swarm movement in time.
Figure C.13. SSA applied to a Levy N.13 function: $F(x,y)=\sin^{2}(3\pi x)+(x-1)^{2}[1+\sin^{2}(3\pi y)]+(y-1)^{2}[1+\sin^{2}(2\pi y)]$. The parameters used are $N=16$, $\lambda=2$ and $\beta=1/8$. The global minimum is $F(x_{\ast},y_{\ast})=0$ at $(x_{\ast},y_{\ast})=(1,1)$. Top Left: landscape of he function. Lower Left: long time behavior. Center and Right: swarm movement in time. Similar to Figure C.12, we can see the provisional minimum converges to the global minimum as the particles are able to accumulate around the global minimum.
Figure C.14. SSA applied to a Drop-Wave function: $F(x,y)=-\frac{1+\cos(12\sqrt{x^{2}+y^{2}})}{0.5(x^{2}+y^{2})+2}$. The parameters used are $N=32$, $\lambda=2$ and $\beta=1/16$. The global minimum is $F(x_{\ast},y_{\ast})=-1$ at $(x_{\ast},y_{\ast})=(0,0)$. Top Left: landscape of he function. Lower Left: long time behavior. Center and Right: swarm movement in time. Top Left: landscape of he function. Lower Left: long time behavior. Center and Right: swarm movement in time. This is a challenging example since the global basin is very small.
Figure C.15. SSA applied to a Schaffer N.2 function: $F(x,y)=0.5+\frac{\sin^{2}(x^{2}-y^{2})-0.5}{[1+0.001(x^{2}+y^{2})]^{2}}$. The parameters used are $N=64$, $\lambda=2$ and $\beta=1/32$. The global minimum is $F(x_{\ast},y_{\ast})=0$ at $(x_{\ast},y_{\ast})=(0,0)$. Top Left: landscape of he function. Lower Left: long time behavior. Center and Right: swarm movement in time. This is also a challenging example since the global basin is very small.
Figure C.16. SSA applied to a Bukin N.6 function: $F(x,y)=100\sqrt{|y-0.01x^{2}|}+0.01|x+10|$. The parameters used are $N=16$, $\lambda=0.25$ and $\beta=1/16$. The global minimum is $F(x_{\ast},y_{\ast})=0$ at $(x_{\ast},y_{\ast})=(-10,1)$. Top Left: landscape of he function. Lower Left: long time behavior. Center and Right: swarm movement in time. This is a very special example: it is observed that the particles can reach the global basin with ease but since all local and global basins are very shallow, samples with Brownian motion easily jump out. As a consequence, the provisional minimum oscillates very wildly.
Figure C.17. SSA applied to a Rastrigin function: $F(x,y)=20+2x^{2}+2y^{2}-\cos(2\pi x)-\cos(2\pi y)$. The parameters used are $N=16$, $\lambda=2$ and $\beta=1/8$. The global minimum is $F(x_{\ast},y_{\ast})=0$ at $(x_{\ast},y_{\ast})=(0,0)$. Top Left: landscape of he function. Lower Left: long time behavior. Center and Right: swarm movement in time. Similar to the 1D Rastrigin case (see Figure 6.11), the provisional minimum quickly decreases to a plateau, and stays there for a long time frame before descending again. The major challenge in this example is that the local minimum has very similar value to the global minimum, and often, when found, trick the swarm to saturates.